< Review of Energy and Work (Chapters 9 and 10) Problem 9.62 34 ot 38> Part A When you ride a bicycle at constant speed, nearly all the energy you expend goes into the work you do against the drag force of the air. Model a cyclist as having cross-section area 0.42 m2 and, because the human body is not aerodynamically shaped, a drag coefficient of 090 What is the cyclists power output while riding at a steady 7.3 m/s (16 mph)? Express your answer to two significant figures and include the appropriate units. P- 180 Submit vious Ans Incorrect: Try Again; 9 attempts remaining Part B Metabolic power is the rate at which your body burns fuel to power your activities For many activities, your body is roughly 25% efficient at converting the chemical energy of food into mechanical energy. What is the cyclists metabolic power while cycling at 7.3 m/s? Express your answer to two significant figures and include the appropriate units. PMValue Units

Could I get some help on this question please?

Also show your work if possible and have a wonderful day!
Answer #1

Solution-

a)Taking the density of air ρ = 1.2 kg/m3

Drag Force = Cd*1/2*ρAV^2

D = 0.9*1/2*1.2*0.42*7.3^2

D = 12.09 N

Power = D*V = 12.09 *7.3 = 88.26 W

b)Metabolic power = 88.26/0.25 = 353.04 W